# Revisiting 360 Depth Estimation with PanoGabor: A New Fusion Perspective

### v1

@@ III-A Distortion-aware PanoGabor Filters (PGF)
We aim to utilize the advanced texture analysis capabilities and inherent frequency domain transformation properties of Gabor filters to address the challenges posed by the large FoV in 360 images. However, due to the severe distortion, directly incorporating the standard Gabor filters into the learning frameworks for monocular 360 depth estimation would lead to unsatisfactory prediction. In contrast, the learnable Gabor filters can dynamically adapt to the distortions but deliver suboptimal performance because the distortion distribution pattern is hard to model in a data-driven way. Inspired by the spherical projection process, we propose to calculate the distortion coefficients for distortion-aware Gabor filter generation. Here, we propose the latitude-aware distortion representation method, which approximates the distortion coefficients as a function of the latitude.
@@
The approximation derives from the equirectangular projection and the geometric properties of a sphere. Due to the spherical representation, the distance covered by a degree of longitude decreases from the equator towards the two poles. For a given latitude $\phi$, the distortion coefficient $c$ can be approximated by the reciprocal of the cosine of the latitude (the proof can be found in the suppl. material):
$c\approx\cos(\phi),$ (1)
the scale factor mirrors the increasing distortion in equirectangular projections towards the poles. However, in practice, we find that this nonlinear cosine representation introduces complex frequency changes making it more difficult to converge stably during training. Although it complies with certain distortion patterns, the instability in the optimization process may offset this advantage. To this end, we introduce a linear distortion representation method:
$c=|\frac{\phi\cdot\pi}{2}|,$ (2)
@@
where $\phi\in[-\frac{\pi}{2},\frac{\pi}{2}]$. As detailed in Algorithm 1, we employ the defined distortion coefficient to control the generation of distortion-aware PanoGabor filters, thereby creating a bank of PanoGabor filters:
$PGF_{i}=$$\{g(x,y,f,\theta_{i},\psi,\sigma)\}_{i=1}^{8},$ (3)
where $g(x,y,f,\theta_{i},\psi,\sigma)$ indicates the Gabor transformation (preliminary definition can be found in the suppl. materials).
Figure 3: Visualization of the features extracted through the vanilla CNN layer and our proposed PanoGabor layer.
PanoGabor Convolution Layer. We further enclose the PanoGabor filters as an independent convolution layer. Given a set of input features $\mathcal{F}\in\mathbb{R}^{B\times C\times H\times W}$, we initially apply a standard convolutional layer to transform these features into channel-height dimension-matched features, denoted as $\mathcal{F^{\prime}}\in\mathbb{R}^{B\times H\times H\times W}$. For each channel, we then calculate a distortion coefficient based on the latitudinal distribution (i.e., $\phi\in[-\frac{\pi}{2},\frac{\pi}{2}]$). Using Algorithm 1, we generate a set of PanoGabor filters for each channel to compensate for the geometric distortion of spherical features. By further making the PanoGabor layer learnable, we enable the PanoGabor filters to become both distortion-aware and data-driven, thereby seamlessly integrating into the subsequent CS-UFM module for joint optimization. As illustrated in Figure 3, the proposed PanoGabor layer demonstrates its capability to reduce the interference of surface texture, thereby effectively extracting ERP features with explicit geometric structures from 360 images.
III-B Spherical Gradient Constraint
@@
We have introduced our PanoGabor approach, which extends the original Gabor filters to be learnable and panorama-specific to achieve the dynamic perception of panoramic features. However, because of the sensitivity to the orientations, PanoGabor does not work well when the original gradient constraints are naively applied. Therefore, we perform a spherical gradient constraint for PanoGabor to assist in the dynamic parameter adjustment process. Specifically, we calculate these gradients on a pixel-level tangent plane. Mathematically, given a tangent sampling [26, 4, 19] function $f_{s}(*)$ over the sphere which can be employed to sample a 3×3 patch for each pixel, the spherical gradients can be defined as follows:
$G_{x}(p)$ $=$ $f_{s}(p)\times S_{x}$ (4)
$G_{y}(p)$ $=$ $f_{s}(p)\times S_{y},$ (5)
where $\phi$ and $\lambda$ represent latitude and longitude of a sphere, respectively; $S_{x},S_{y}$ represent the horizontal/vertical part of the Sobel operator.
III-C Channel-wise and Spatial-wise Unidirectional Fusion Module (CS-UFM)
@@
The fusion module is designed to integrate two sets of ERP features, addressing reintroduced distortion, inconsistencies in cubemap boundaries, and pixel loss issues. This module first concatenates the two groups of input ERP features and uses a residual block to achieve an initial refinement. In particular, the block comprises a 1x1 convolution that compresses the channels to match the features’ height, facilitating the application of the PanoGabor layer. Following the 1×1 convolution layer, a 3x3 PanoGabor layer is used to process the features, aiming for distortion awareness and depth cues enhancement.
@@
Figure 4: The proposed CS-UFM module.
@@
Datasets Backbone Method Pub’Year Params (M) FLOPs $(\bar{G})$ Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$ $\delta_{2}(\%)\uparrow$ $\delta_{3}(\%)\uparrow$
Transformer EGFormer ICCV’23 15.39 66.21 0.1473 0.1517 0.6025 81.58 93.90 97.35
PanoFormer ECCV’22 20.38 81.09 0.1051 0.0966 0.4929 89.08 96.23 98.31
BiFuse CVPR’20 56.01 199.58 0.1126 0.0992 0.5027 88.00 96.13 98.47
BiFuse++ TPAMI’22 52.49 87.48 0.1123 0.0915 0.4853 88.12 96.56 98.69
M3D ResNet-34 UniFuse RAL’21 50.48 96.52 0.1144 0.0936 0.4835 87.85 96.59 98.73
OmniFusion CVPR’ 22 42.46 142.29 0.1161 0.1007 0.4931 87.72 96.15 98.44
HRDFuse CVPR’23 46.31 80.87 0.1172 0.0971 0.5025 86.74 96.17 98.49
Elite360D CVPR’24 25.54 65.29 0.1115 0.0914 0.4875 88.15 96.46 98.74
Ours - 30.75 75.67 0.1067 0.0858 0.4624 89.06 96.75 98.72
Transformer EGFormer ICCV’23 15.39 66.21 0.1528 0.1408 0.4974 81.85 93.38 97.36
PanoFormer ECCV’22 20.38 81.09 0.1122 0.0786 0.3945 88.74 95.84 98.59
OmniFusion CVPR’22 42.46 142.29 0.1154 0.0775 0.3809 86.74 96.03 98.71
S2D3D ResNet-34 UniFuse RAL’21 50.48 96.52 0.1124 0.0709 0.3555 87.06 97.04 98.99
Elite360D CVPR’24 25.51 65.28 0.1182 0.0728 0.3756 88.72 96.84 98.92
Ours - 30.75 75.67 0.1029 0.0599 0.3228 89.13 96.87 99.06
Transformer EGFormer ICCV’23 15.39 66.21 0.2205 0.4509 0.6841 79.79 90.71 94.55
PanoFormer ECCV’22 20.38 81.09 0.2549 0.4949 0.7937 74.70 89.15 93.97
Struct3D BiFuse CVPR’ 20 56.01 199.58 0.1573 0.2455 0.5213 85.91 94.00 96.72
ResNet-34 UniFuse RAL’ 21 50.48 96.52 0.1506 0.2319 0.5016 85.42 93.99 96.76
Elite360D CVPR’24 25.51 65.28 0.1480 0.2215 0.4961 87.41 94.34 96.66
Ours - 30.75 75.67 0.1401 0.2499 0.4394 87.78 94.49 96.82
III-D Loss Function
We employ the widely used reverse Huber Loss, and the proposed Spherical Gradient Loss to form our objective function. The standard loss functions of the former can be written as follows:
$\mathcal{L}_{\beta}(d,\hat{d})=\frac{1}{N}\sum\limits_{i\in N}\left\{\begin{array}[]{ll}|d_{i}-\hat{d_{i}}|&|d_{i}-\hat{d_{i}}|\leq\theta\\ \frac{(d_{i}-\hat{d_{i}})^{2}+\delta^{2}}{2\delta}&\text{ otherwise },\end{array}\right.$ (8)
where $d$/$\hat{d}$ is the ground-truth/predicted-value, $\theta$=0.2. For the Spherical Gradient Loss, we represent it as follows:
[…]
Ultimately, we formulate the objective function as follows:
[…]
where $\eta=0.5$; $G$: ground truth; $D$: predicted depth map.
@@ IV Experiment
We conduct all experiments on a single NVIDIA GeForce RTX 3090 GPU, implementing our approach using the PyTorch framework. We use the Adam optimizer [29] with an initialized learning rate of $1\times 10^{-4}$ to train our model. We keep the default settings for other hyper-parameters of the optimizer. During the training period, we use a batch size of 8 and employ left-right flipping, panoramic horizontal rotation, and luminance change for data augmentation, as previously proposed in works [2, 4]. The backbones we employed are all pre-trained on ImageNet-1K.
Datasets. For experimental validation, two real-world datasets (Stanford2D3D [8], and Matterport3D [9]) and a synthetic dataset (Structured3D [10]). We adopt a resolution of 1024×512 (width×height, 3 channels) of the single input panorama for all the datasets.

### v4

@@ III-B Distortion-aware PanoGabor Filters (PGF)
We aim to utilize the advanced texture analysis capabilities and inherent frequency domain transformation properties of Gabor filters to address the challenges posed by the large FoV in 360 images. However, due to the severe distortion, directly incorporating the standard Gabor filters into the learning frameworks for monocular 360 depth estimation would lead to unsatisfactory prediction. In contrast, the learnable Gabor filters can dynamically adapt to the distortions but deliver suboptimal performance because the distortion distribution pattern is hard to model in a data-driven way. Inspired by the spherical projection process, we propose calculating the distortion coefficients for distortion-aware Gabor filter generation. Here, we propose the latitude-aware distortion representation method, which approximates the distortion coefficients as a function of the latitude.
@@
The approximation derives from the equirectangular projection and the geometric properties of a sphere. Due to the spherical representation, the distance covered by a degree of longitude decreases from the equator towards the two poles. For latitudes $\phi\in(-\frac{\pi}{2},\frac{\pi}{2})$, the distortion coefficient $d_{c}$ can be approximated by the reciprocal of the cosine of the latitude (the proof can be found in Sec. III-A):
$K(\phi)=\frac{1}{\cos(\phi)}-1,$ (8)
@@
where the range of $d_{c}$ is $[0,+\infty)$. The scale factor mirrors the increasing distortion in equirectangular projections towards the poles. Subsequently, as detailed in Algorithm 1, we employ the defined distortion coefficient to control the generation of distortion-aware PanoGabor filters, thereby creating a bank of PanoGabor filters.
Figure 4: Visualization of the features extracted through the vanilla CNN layer and our proposed PanoGabor layer. Compared to the vanilla CNN layer, our PanoGabor layer is more robust regarding surface textures and complex illumination variations, which is crucial for accurate indoor depth estimation.
PanoGabor Convolution Layer. We further enclose the PanoGabor filters as an independent convolution layer. Given a set of input features $\mathcal{F}\in\mathbb{R}^{B\times C\times H\times W}$, we initially apply a standard convolutional layer to transform these features into channel-height dimension-matched features, denoted as $\mathcal{F^{\prime}}\in\mathbb{R}^{B\times H(C_{in})\times H\times W}$. Firstly, we calculate the distortion coefficient for each channel. Let $C=[0,1,\dots,C_{in}-1]$ be the latitude grid indices, the latitude value can be calculated as follows:
$\phi_{C}=\left\{(\frac{c_{i}}{C_{in}}-0.5)\pi|c_{i}\in C\right\}.$ (9)
We then apply Eq. 8 to calculate the distortion coefficients $K(\phi_{C})\in\mathbb{R}^{C_{in}}$ for all channels. Subsequently, we use Algorithm 1 to generate a set of PanoGabor filters. These filters are then employed to perform the convolution operation, compensating for the geometric distortions in spherical features. By making the PanoGabor layer learnable, the filters become distortion-aware and data-driven. This allows the PanoGabor layer to integrate seamlessly with the subsequent CS-UFM module, facilitating joint optimization. As illustrated in Fig. 4, compared to the vanilla CNN layer, our PanoGabor layer demonstrates enhanced robustness to interference from surface textures and complex illumination variations in indoor scenes, which is crucial for accurate indoor depth estimation. Furthermore, our method remains unaffected by distortions, thereby effectively extracting ERP features with explicit geometric structures from 360 images.
Figure 5: Illustration of the spherical tangent sampling. a represent that the pixels are projected from the pixel coordinates to the spherical coordinates to sampling tangent points. b indicates that the calculated sampling points are projected back to the pixel coordinates.
III-C Spherical Gradient Constraint
@@
We have introduced our PanoGabor approach, which extends the original Gabor filters to be learnable and panorama-specific to achieve the dynamic perception of panoramic features. However, because of the sensitivity to the orientations, PanoGabor does not work well when the original gradient constraints are naively applied. Therefore, we perform a spherical gradient constraint for PanoGabor to assist in the dynamic parameter adjustment. Specifically, we calculate gradients on a pixel-level tangent plane, following the methodology outlined in previous works [36, 37].
Create sampling mesh grid. Given a depth map $D\in\mathbb{R}^{H\times W}$, we first define the sampling grid. For each pixel in the depth map, we sample a 3×3 tangent patch. Let $X=[-1,0,1]$ and $Y=[-1,0,1]$ represent the sampling kernels for the $x$- and $y$-directions, respectively. We define the 2D grid mesh $M$ (Fig. 5left) as:
$M=\left\{(M_{x}(x_{i}),M_{y}(y_{j}))\mid x_{i}\in X,y_{j}\in Y\right\},$ (10)
where $M_{x}(x_{i})$ and $M_{y}(y_{j})$ represent the corresponding coordinates in the tangent plane for the sampling positions:
$M_{x}(x_{i})$ $=\tan(x_{i}\cdot\Delta\lambda),$ (11)
$M_{y}(y_{j})$ $=\frac{\tan(y_{j}\cdot\Delta\phi)}{\cos(y_{j}\cdot\Delta\lambda)}.$
Here, $\Delta\lambda=\frac{2\pi}{W}$ and $\Delta\phi=\frac{\pi}{H}$ denote the sampling step sizes in the longitude $\lambda$ and latitude $\phi$ directions, respectively. This ensures that the sampling grid appropriately covers the spherical surface based on the pixel indices.
Pixel coordinates to spherical coordinates (Fig. 5a). For each pixel located at row $h$ and column $w$ of the depth map, we compute the corresponding latitude and longitude values in spherical coordinates. The latitude $\phi_{h}$ for row $h$ is calculated as:
$\phi_{h}=\left(\frac{h}{H}-0.5\right)\pi,$ (12)
and the longitude $\lambda_{w}$ for column $w$ is given by:
$\lambda_{w}=\left(\frac{w}{W}-0.5\right)2\pi.$ (13)
These latitude and longitude values correspond to the pixel’s position on the spherical surface, with the normalization accounting for the pixel dimensions in the depth map.
Calculate radial distance and angular distortion. For each sampled kernel point at position $(x_{i},y_{j})$, we compute the radial distance $\rho$, which is the Euclidean distance between the corresponding coordinates in the tangent plane:
$\rho(x_{i},y_{j})=\sqrt{M_{x}(x_{i})^{2}+M_{y}(y_{j})^{2}}.$ (14)
The radial distance is then used to calculate the angular distortion $\nu$, which accounts for the curvature of the spherical surface:
$\nu(x_{i},y_{j})=\arctan(\rho(x_{i},y_{j})).$ (15)
This angular distortion is crucial for compensating for the non-uniformity introduced by the spherical geometry.
Latitude and longitude of tangent grid points. Using the radial distance $\rho(x_{i},y_{j})$ and angular distortion $\nu(x_{i},y_{j})$, we compute the latitude and longitude for each sampling point in the tangent patch. The latitude $\phi(h,x_{i},y_{j})$ for the pixel $(h,w)$ and the kernel position $(x_{i},y_{j})$ is computed as:
$\phi(h,x_{i},y_{j})=\arcsin\Big{(}\cos(\nu(x_{i},y_{j}))\sin(\phi_{h})$ (16)
$+\frac{M_{y}(y_{j})\sin(\nu(x_{i},y_{j}))\cos(\phi_{h})}{\rho(x_{i},y_{j})}\Big{)}.$
Similarly, the longitude $\lambda(h,x_{i},y_{j})$ is calculated as:
$\lambda(h,x_{i},y_{j})=\lambda_{1}(h,x_{i},y_{j})+\lambda_{w},$ (17)
where $\lambda_{1}$ is given by:
$\lambda_{1}(h,x_{i},y_{j})=\arctan\left(\frac{M_{x}(x_{i})\sin(\nu(x_{i},y_{j}))}{\Delta}\right),$ (18)
and $\Delta$ is computed as:
$\Delta=\rho(x_{i},y_{j})\cos(\phi_{h})\cos(\nu(x_{i},y_{j}))$ (19)
$-M_{y}(y_{j})\sin(\phi_{h})\sin(\nu(x_{i},y_{j})).$
Projection to pixel coordinates. Finally, we project the calculated latitude and longitude values back onto the image pixel coordinates (Fig. 5b). The new row and column positions $h^{\prime}$ and $w^{\prime}$ are given by:
$h^{\prime}$ $=\frac{\phi(h,x_{i},y_{j})H}{\pi}+0.5,$ (20)
$w^{\prime}$ $=\left(\frac{\lambda(h,x_{i},y_{j})}{2\pi}+0.5\right)W\mod W,$
where the MOD operation appropriately handles the seamless nature of the left and right boundaries in panoramic images, ensuring that when the sampling range exceeds the boundaries, it is wrapped around to the opposite side. With these projected positions, we sample the corresponding depth values from the depth map. The function $f_{s}(*)$ formalizes this sampling process, which allows us to efficiently compute the gradients on the spherical surface.
Gradient calculation. The gradients are then computed by applying the Sobel operator in the $x$- and $y$-directions. The horizontal and vertical gradients $G_{x}(D)$ and $G_{y}(D)$ are given by:
$G_{x}(D)$ $=\frac{1}{HW}\sum_{i=0}^{HW-1}f_{s}(D_{i})\otimes S_{x},$ (21)
$G_{y}(D)$ $=\frac{1}{HW}\sum_{i=0}^{HW-1}f_{s}(D_{i})\otimes S_{y},$
where $S_{x}$ and $S_{y}$ represent the horizontal and vertical components of the Sobel operator, respectively. These gradients are computed over the entire depth map $D$, where $HW$ is the total number of pixels. The use of the Sobel operator ensures that the gradients capture both the local spatial variations and the curvature effects on the spherical surface.
Figure 6: The proposed CS-UFM module.
III-D Channel-wise and Spatial-wise Unidirectional Fusion Module (CS-UFM)
@@
The fusion module is designed to integrate two sets of ERP features, addressing reintroduced distortion, inconsistencies in cubemap boundaries, and pixel loss issues. This module (as illustrated in Fig. 6) first concatenates the two groups of input ERP features and uses a residual block to achieve an initial refinement. In particular, the block comprises a 1x1 convolution that compresses the channels to match the features’ height, facilitating the application of the PanoGabor layer. Following the 1×1 convolution layer, a 3x3 PanoGabor layer is used to process the features, aiming for distortion awareness and depth cues enhancement.
@@
III-E Loss Function
We propose an objective function that combines reverse huber loss and spherical gradient loss to jointly optimize pixel-wise depth accuracy and spatial gradient consistency of the predicted depth maps. The Reverse Huber Loss is defined as:
$\mathcal{L}_{\beta}(D,G)=\frac{1}{N}\sum_{i\in N}\begin{cases}|D_{i}-G_{i}|&\text{if }|D_{i}-G_{i}|\leq\theta,\\ \frac{(D_{i}-G_{i})^{2}+\delta^{2}}{2\delta}&\text{otherwise},\end{cases}$ (22)
where $D_{i}$ and $G_{i}$ are the predicted and ground truth depth values at the $i$-th pixel, respectively, and $N$ is the total number of pixels in the depth map. The threshold $\theta=0.2$ determines when the error is treated linearly or quadratically, while $\delta$ is a small constant that smooths the quadratic penalty for large errors. This loss function applies a linear penalty for small errors and a smooth quadratic penalty for large errors, providing robustness against noise and outliers.
The spherical gradient loss is defined as:
$\mathcal{L}_{g}=|G_{x}(G)-G_{x}(D)|+|G_{y}(G)-G_{y}(D)|,$ (23)
where $G$ and $D$ represent the ground truth and predicted depth maps, respectively, and $G_{x}$ and $G_{y}$ denote the gradients of these depth maps along the horizontal and vertical directions. This loss encourages the predicted depth map to maintain spatial gradient consistency with the ground truth, helping preserve smoothness and structure.
Finally, the overall objective function is:
$\mathcal{L}=\mathcal{L}_{\beta}(D,G)+\eta\mathcal{L}_{g},$ (24)
where $\eta=0.5$ is a hyperparameter that balances the contributions of the two loss terms. By jointly optimizing these two losses, the model learns to predict accurate depth values while maintaining spatial structure and smoothness in the predicted depth maps.
[b] Datasets Backbone Method Pub’Year $(\bar{G})$ Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$ $\delta_{2}(\%)\uparrow$ $\delta_{3}(\%)\uparrow$ Transformer EGFormer ICCV’23 0.1473 0.1517 0.6025 81.58 93.90 97.35 PanoFormer ECCV’22 0.1051 0.0966 0.4929 89.08 96.23 98.31 BiFuse CVPR’20 0.1126 0.0992 0.5027 88.00 96.13 98.47 BiFuse++ TPAMI’22 0.1123 0.0915 0.4853 88.12 96.56 98.69 M3D UniFuse RAL’21 0.1144 0.0936 0.4835 87.85 96.59 98.73 ResNet-34 OmniFusion CVPR’ 22 0.1161 0.1007 0.4931 87.72 96.15 98.44 HRDFuse CVPR’23 0.1172 0.0971 0.5025 86.74 96.17 98.49 Elite360D CVPR’24 0.1115 0.0914 0.4875 88.15 96.46 98.74 Ours - 0.1067 0.0858 0.4624 89.06 96.75 98.72 Transformer EGFormer ICCV’23 0.1528 0.1408 0.4974 81.85 93.38 97.36 PanoFormer ECCV’22 0.1122 0.0786 0.3945 88.74 95.84 98.59 OmniFusion CVPR’22 0.1154 0.0775 0.3809 86.74 96.03 98.71 S2D3D UniFuse RAL’21 0.1124 0.0709 0.3555 87.06 97.04 98.99 ResNet-34 Elite360D CVPR’24 0.1182 0.0728 0.3756 88.72 96.84 98.92 Ours - 0.1029 0.0599 0.3228 89.13 96.87 99.06 Transformer EGFormer ICCV’23 0.2205 0.4509 0.6841 79.79 90.71 94.55 PanoFormer ECCV’22 0.2549 0.4949 0.7937 74.70 89.15 93.97 Struct3D BiFuse CVPR’ 20 0.1573 0.2455 0.5213 85.91 94.00 96.72 UniFuse RAL’ 21 0.1506 0.2319 0.5016 85.42 93.99 96.76 ResNet-34 Elite360D CVPR’24 0.1480 0.2215 0.4961 87.41 94.34 96.66 Ours - 0.1401 0.2499 0.4394 87.78 94.49 96.82
@@ III-E Loss Function
  • The bold values indicate the best performance.
TABLE II: Complexity comparison with the SOTA methods.
Backbone Method Params (M) FLOPs
Transformer EGFormer 15.39 66.21
PanoFormer 20.38 81.09
BiFuse 56.01 199.58
BiFuse++ 52.49 87.48
ResNet-34 UniFuse 50.48 96.52
OmniFusion 42.46 142.29
HRDFuse 46.31 80.87
Elite360D 25.54 65.29
Ours 30.75 75.67
@@ IV Experiment
We conduct all experiments on a single NVIDIA GeForce RTX 3090 GPU, implementing our approach using the PyTorch framework. We use the Adam optimizer [39] with an initialized learning rate of $1\times 10^{-4}$ to train our model. We keep the default settings for other optimizer hyper-parameters. During the training period, we use a batch size of 8 and employ left-right flipping, panoramic horizontal rotation, and luminance change for data augmentation, as previously proposed in works [2, 4]. The backbones we employed are all pre-trained on ImageNet-1K.
Dataset. Two real-world datasets for experimental validation: Stanford 2D-3D [12], and Matterport3D [40]. The depth maps in Stanford 2D-3D are constructed from reconstructed 3D models, and some depths are inaccurate. Besides, the bottom and top regions of the RGB images are missing. We follow the official splits and get 1,040 RGB-D pairs for training and 343 for testing. The Matterport3D dataset provides raw depth whose top and bottom depths are missed. For the RGB panoramas in Matterport3D, we follow the Unifuse method [41] to fill the missing regions. Finally, we get 7829 pairs for training and 2014 pairs for testing.
One virtue dataset: Structrued3D [14]. The virtue dataset was first proposed to explore the layout estimation task. Recently, some researchers have preferred to evaluate the depth estimation models on the dataset for its high-quality annotations. We follow the official splits that we employ the frame scene_0000 to frame scene_02999 for training and scene_03250 to scene_03499 for testing. We especially follow the official guidelines and remove invalid cases.
For all the datasets, we adopt a resolution of 1024×512 (width×height, 3 channels) of the single input panorama.
Figure 8: Visualization of the point cloud generated with the ground truth (GT) and the predicted depth map (Ours and UniFuse). The areas marked in red rectangles indicate that our approach gains superior performance (dense and high-quality point cloud).